B=c/1-u Solve for u.
Answer by texttutoring(324)

B=c/1-u

First, multiply both sides by the denominator (1-u) to move it over to the top of the left side.

B(1-u) = c

Now divide both sides by B:

1-u = c/B

Now add u to both sides:
1 = c/B + u

Now subtract c/B from both sides:
1-c/B = u